OpenStudy (jaylelile):

If 14/3=x/y, then 14/x= ?

OpenStudy (anonymous):

I think it's one but I'm not sure.

OpenStudy (anonymous):

Did you mean: \[\frac{ 14 }{ 3 }=\frac{ x }{ y }\] then \[\frac{ 14 }{ x }=?\]

OpenStudy (jaylelile):

yes, that is what I mean.

OpenStudy (jaylelile):

My options are:

OpenStudy (anonymous):

Ok so then it's D \[\frac{ 3 }{ y }\]

OpenStudy (jaylelile):

Thank you!